# Vascular Ageing and Smoking Habit Prediction via a Low-Cost Single-Lead ECG Module

### v2

@@ Abstract
This paper presents a novel low-cost method to predict: i) the vascular age of an apparently healthy young person, ii) whether or not a person smokes, using only the lead-I of the electrocardiogram (ECG). We begin by collecting (lead-I) ECG data from 42 subjects (male, female, smoker, non-smoker) aged 18 to 30 years, using our custom-built low-cost single-lead ECG module, and anthropometric data, e.g., body mass index, smoking status, blood pressure etc. Under our proposed method, we first pre-process our dataset by denoising the ECG traces, followed by baseline drift removal, followed by z-score normalization. Next, we divide ECG traces into overlapping segments of five-second duration, which leads to a 145-fold increase in the size of the dataset. We then feed our dataset to a number of machine learning models, a 1D convolutional neural network, a multi-layer perceptron (MLP), and ResNet18 transfer learning model. For vascular ageing prediction, Random Forest method outperforms all other methods with an $R^{2}$ score of 0.99, and mean squared error of 0.07. For the binary classification problem (smoker or non-smoker), XGBoost method stands out with an accuracy of 96.5%. Finally, for the 4-class classification problem that aims to differentiate between male smoker, female smoker, male non-smoker, and female non-smoker, MLP method achieves the best accuracy of 97.5%. This work is aligned with the sustainable development goals of the United Nations which aim to provide low-cost but quality healthcare sensing solutions to the unprivileged.
@@ I Introduction
As discussed above, vascular age and smoking both result in alterations in the ECG [4], [5]. This motivates us to study the feasibility of using ECG as a clinical tool to assess the impact of vascular age and smoking on cardiovascular health. This objective is inline with the broad research efforts that aim to use ECG as a diagnostic tool with the ultimate objective of bringing down the annual CVD-based deaths by early detection of CVDs [6]. A reduction in proliferation of CVDs will imply a reduction in health burden of countries in lieu of health insurance costs. However, note that manual assessment of the ECG to look for possible irregularities in the heart rhythm due to vascular ageing and smoking is time-consuming and error-prone. Hence, artificial intelligence (AI)-based methods which could automatically and reliably predict the vascular age [7] and smoking status of a person (smoker or non-smoker) from the (single-lead) ECG data are the need of the hour. This is precisely the scope of this work.
@@
Contributions.
  • We design a low-cost single-lead ECG sensor module (with a price tag of 10 USD), and utilize it to collect (lead-I) ECG data from 42 subjects (male, female, smoker, non-smoker) aged 18 to 30 years.
  • We feed our custom dataset to various machine, deep and transfer learning methods, which reliably predict the vascular age of a person using a low-cost single-lead (lead-I) ECG sensor module only. This is a giant leap compared to the related work that either utilizes the 12-lead ECG [5] or utilizes invasive methods (e.g., central aortic blood pressure monitoring) for the same task.
  • We also pass our dataset to a number of machine and deep learning methods, which reliably differentiate between a smoker and a non-smoker solely based on the single-lead ECG data. To the best of our knowledge, this work is the first that aims to infer the smoking status of a person from his/her single-lead ECG data.
Finally, note that this work is also aligned with the sustainable development goals (3 and 10) of the United Nations which aim to provide low-cost but quality healthcare sensing solutions to the unprivileged populations (e.g., people in developing countries and people in remote areas).
Outline. Section II provides a summary of the selected related work. Section III describes the proposed methodology as well as the hardware architecture of our custom low-cost single-lead ECG module. Section IV discusses the data collection process for our custom ECG dataset, the data pre-processing steps, and feature extraction process. Section V enumerates the ML, DL and TL models we have implemented. Section VI describes selected results for both problems (vascular ageing prediction and smoking habit prediction). Finally, Section VII concludes the paper.
@@ II Related work
To date, a multitude of methods have been investigated, which either utilize the 12-lead ECG [5] or PPG sensors [9].
@@
Statistical methods: [10] develops a Bayesian statistical model that predicts an individual’s heart age based on his/her resting 5 min 12-lead ECG. They evaluated their model on 776 healthy individuals, 221 with cardiovascular risk factors, 441 with cardiac disease, and a small group of highly endurance-trained athletes. Their model predicts a higher heart age for the subjects with risk factors, with heart diseases, and athletes. Motivated by the situations where cfPWV cannot be measured, [11] computes the estimated pulse wave velocity (ePWV) from the age and brachial blood pressure. They measure cfPWV in two-hundred and fifty-two adults, and compare it with each participant’s ePWV, in order to correlate the ePWV with cfPWV (which is an established measure of vascular aging).
@@
Machine learning based methods: Recently, machine learning methods for vascular ageing prediction have gained widespread attention (see the survey article [7] and references therein for a quick review of the development in the field). Due to space constraints, we discuss selected works below. [12] utilizes a 13-years MESA dataset of 6459 atherosclerotic CVD-free participants and demonstrates that their proposed support vector machine (SVM)-based CVD risk calculator outperforms the classical PCE-based CVD risk calculator. [13] utilizes a public dataset of video PPG signals acquired from smartphones, does the feature extraction, and trains a number of machine and deep learning methods in order to predict healthy vascular ageing. Authors of [14] take a critical approach and study the bias problem in ML-based vascular age prediction methods. They argue that overfitting and lack of large datasets (with diverse ethnic populations) might be the reasons behind the fact that ML methods yield inaccurate results during the clinical trials. Similarly, [15] argues that it is possible to accurately predict patients’ Bayesian 5-min ECG Heart ages from their standard, resting 10-s 12-lead ECGs using standard regression models only, without relying upon the AI techniques. They utilize a public dataset with a total of 2,771 subjects (1682 healthy volunteers, 305 with cardiovascular risk factors, 784 with cardiovascular disease).
@@
Deep learning based methods: In [16], authors train a residual neural network on the 12-lead ECG data from the CODE study cohort (1,558,415 patients), and report that the patients with a heart age gap of more than 8 years have a higher mortality rate. [17] trains a convolutional neural network (CNN) on 12-lead ECG data of 17,042 subjects, and reports that the age predicted by their model matches well with the chronological age of young population, but not for the elderly (with age 6060$\ge$ 6060 years). [18] trains the ECG12Net model (basically, a CNN with 82 convolutional layers) on 71,741 subjects (20 to 80 years old), and reports that the high ECG-age is not only correlated with mortality but with several CVD-related outcomes as well. Authors of [19] collect PPG data from 757 participants, utilize a Gaussian mixture model to decompose the PPG signal into incident and reflected waves for feature extraction, and eventually train a shallow artificial neural network for vascular ageing prediction. [20] trains a CNN on 10-second 12-lead ECG data and CVD-related metadata from 499,727 subjects, with the aim to learn the genetic basis of cardiovascular ageing, in order to support future personalised/precision medicine (when genomic information is readily available). In [21], researchers train their CNNs using 10-second samples of 12-lead ECG signals from 499,727 patients in order to predict their biological age and gender. Authors of [22] train a CNN on over 700,000 ECGs from the Mayo Clinic, U.S.A, in order to later predict the ECG-age of 4,542 participants in two cities in Russia. Authors report that $\delta$-age (i.e., ECG-derived artificial intelligence-estimated age minus chronological age) biomarker is strongly correlated with the CVD risk factors, e.g., blood pressure, BMI, total cholesterol and smoking. [23] trains a CNN on the 12-lead ECG data of 774,783 unique subjects, and conclude that the patients with peripheral microvascular endothelial dysfunction (an index of vascular aging) and accelerated physiologic aging are at a greater risk for CVDs.
@@
Low-cost single-lead ECG module design: Since this work utilizes the AD8232 chip (a low-cost single-lead ECG sensor with a price tag of 5 USD), it is imperative to summarize the works that design low-cost single-lead ECG module with the unanimous aim to provide affordable sensing solutions for remote health monitoring. In [27],[28], authors consider a AD8232 chip-based low-cost single-lead ECG module for atrial fibrillation detection (a common cardiac arrhythmia associated with stroke risk). [29] presents a cost-effective method for detecting driver drowsiness, utilizing heart rate variability data from the ECG sensor. [30], [31], [32] all focus on the design of an internet-of-things (IoT)-based low-cost ECG system which reports various biomarkers (e.g., pulse rate, SpO2, body temperature, etc) of remote Covid-19 patients to the doctor, and could detect various cardiac abnormalities such as premature ventricular contraction, premature atrial complex, etc. [33] aims to estimate blood glucose level in a subject using a non-invasive wearable ECG device, using features such as T-wave, ST-segment, etc.
@@ III-A The Proposed Framework
Fig. 1 provides a detailed overview of our proposed method. As can be seen, we utilize our custom low-cost single-lead ECG sensor module in order to construct a labelled dataset. We then do pre-processing of the noisy ECG data, followed by segmentation, followed by feature extraction. Eventually, we pass a total of 21 features (13 ECG-based features and 8 other features) to a number of ML algorithms, a 1D CNN model, and the ResNet18 model which do vascular ageing prediction and smoking habit prediction.
Figure 1: Our proposed method for vascular ageing and smoking habit prediction.
@@ III-B The Low-cost Single-Lead ECG Module
We utilize the AD8232 chip which is a low-cost single-lead ECG sensor (with a price tag of 5 USD). This ECG sensor requires a DC voltage in the range of 2-3.5 volts, and comes with three electrodes which enable it to capture the tiny voltages generated by the human heart, at a rate of 100 Hz. For systematic data acquisition, the AD8232 ECG sensor is interfaced with the ESP32-WROOM micro-controller board (developer version). Further, a 3.7 volts lithium-ion battery is used to power the ESP32 board, the ECG sensor, and TP4056 module (basically, a battery management system). This way, we realize a very cost-effective single-lead ECG sensor module (with a price tag of 10 USD). We connect and solder all the aforementioned hardware components on a veroboard, and then place it in an acrylic body designed via the Solidworks tool. We program the ESP32-WROOM and AD8232 ECG sensor using the Arduino IDE with the objective to read and log the ECG waveform for the purpose of our custom dataset construction. Fig. 2 shows the complete schematic of our custom low-cost single-lead sensor ECG Module.
Figure 2: Our custom single-lead ECG sensor module.
@@ IV-A Data Collection
We utilized our custom module to collect single-lead ECG data from 42 subjects (26 male, 16 female) of age 18 years to 30 years, on campus [^1] This work was approved by the ethical institutional review board (EIRB) of Information Technology University (ITU), Lahore, Pakistan.. Each subject contributed to 3 minutes worth of data, leading to a dataset having 126 minutes (7560 seconds) of single-lead ECG data. Among the male subjects, 16 were smokers, while 10 were non-smokers. Similarly, 4 female subjects were smokers, while 12 were non-smokers. For the ground truth for the vascular ageing prediction problem, we utilized chronological age as a proxy for the vascular age (which is a reasonable assumption keeping in mind that we collected data from apparently healthy young subjects only) [15],[13]. In addition to acquiring the ECG traces of the subjects, we also recorded the relevant metadata such as age, height, weight, heart disease history in the family, smoking habits, sleep duration, blood pressure (systolic and diastolic), body mass index (BMI), and the resting heart rate.

### v3

# Evaluation of a Low-Cost Single-Lead ECG Module for Vascular Ageing Prediction and Studying Smoking-induced Changes in ECG

@@ Abstract
Vascular age is traditionally measured using invasive methods or through 12-lead electrocardiogram (ECG). This paper utilizes a low-cost single-lead (lead-I) ECG module to predict the vascular age of an apparently healthy young person. In addition, we also study the impact of smoking on ECG traces of the light-but-habitual smokers. We begin by collecting (lead-I) ECG data from 42 apparently healthy subjects (smokers and non-smokers) aged 18 to 30 years, using our custom-built low-cost single-lead ECG module, and anthropometric data, e.g., body mass index, smoking status, blood pressure, etc. Under our proposed method, we first pre-process our dataset by denoising the ECG traces, followed by baseline drift removal, followed by z-score normalization. Next, we create another dataset by dividing the ECG traces into overlapping segments of five-second duration. We then feed both segmented and unsegmented datasets to a number of machine learning models, a 1D convolutional neural network, and ResNet18 model, for vascular ageing prediction. We also do transfer learning whereby we pre-train our models on a public PPG dataset, and later, fine-tune and evaluate them on our unsegmented ECG dataset. The random forest model outperforms all other models and previous works by achieving a mean squared error (MSE) of 0.07 and coefficient of determination $R^{2}$ of 0.99, MSE of 3.56 and $R^{2}$ of 0.26, MSE of 0.99 and $R^{2}$ of 0.87, for segmented ECG dataset, for unsegmented ECG dataset, and for transfer learning scenario, respectively. Finally, we utilize the explainable AI framework to identify those ECG features that get affected due to smoking. This work is aligned with the sustainable development goals 3 and 10 of the United Nations which aim to provide low-cost but quality healthcare solutions to the unprivileged. This work also finds its applications in the broad domain of forensic science.
@@ I Introduction
Acronym Definition
ECG Electrocardiogram
CVD Cardiovascular Disease
UN United Nations
SDG Sustainable Development Goal
WHO World Health Organization
AI Artificial Intelligence
ML Machine Learning
DL Deep Learning
TL Transfer Learning
xAI Explainable AI
PPG Photoplethysmography
LR Linear Regression
RR Ridge Regression
DT Decision Tree
RF Random Forest
CNN Convolutional Neural Network
MSE Mean Squared Error
R2 Coefficient of Determination
Table I: List of Acronyms
@@
As discussed above, vascular age and smoking both result in alterations in the ECG [4], [5]. This motivates us to study the feasibility of using ECG as a clinical tool to assess the impact of vascular age and smoking on cardiovascular health. This objective is inline with the broad research efforts that aim to use ECG as a diagnostic tool with the ultimate objective of bringing down the annual CVD-based deaths by early detection of CVDs [6]. A reduction in proliferation of CVDs will imply a reduction in health burden of countries in lieu of health insurance costs. However, note that manual assessment of the ECG to look for possible irregularities in the heart rhythm due to vascular ageing and smoking is time-consuming and error-prone. Hence, artificial intelligence (AI)-based methods which could automatically and reliably predict the vascular age [7] of a person (smoker or non-smoker) from the (single-lead) ECG data are the need of the hour. This is precisely the scope of this work.
@@
Contributions. The main contributions of this paper are three-fold:
  • Data acquisition: We design a low-cost single-lead ECG sensor module (with a price tag of 10 USD), and utilize it to collect (lead-I) ECG data from 42 apparently healthy subjects (male, female, smoker, non-smoker) aged 18 to 30 years. We also segment our dataset in order to get another large dataset that consists of 6,131 data points.
  • Vascular age estimation: We feed the two custom datasets (segmented and unsegmented) to various machine, deep and transfer learning methods, which predict the vascular age of a person using a low-cost single-lead (lead-I) ECG sensor module only. We also do transfer learning whereby we pre-train our models on a public PPG dataset, and later, fine-tune and evaluate them on our unsegmented ECG dataset.
  • Study of smoking-induced changes in single-lead ECG: We also do explainable AI-based feature analysis on the two custom datasets (segmented and unsegmented) in order to identify those features that get affected due to smoking.
We note that our work advances the state-of-the-art that mostly utilizes either the 12-lead ECG [5] or invasive methods (e.g., central aortic blood pressure monitoring) for vascular age estimation. Further, this work is also aligned with the sustainable development goals (3 and 10) of the United Nations which aim to provide low-cost but quality healthcare sensing solutions to the unprivileged populations (e.g., people in developing countries and people in remote areas). In addition, this work also finds its applications in the broad domain of forensic science.
Outline. Section II provides a summary of the selected related work. Section III describes the proposed methodology as well as the hardware architecture of our custom low-cost single-lead ECG module. Section IV discusses the data collection process for our custom ECG dataset, the data pre-processing steps, and feature extraction process. Section V enumerates the ML, DL and TL models we have implemented. Section VI describes selected results for vascular ageing prediction and smoking-induced changes in ECG. Finally, Section VII concludes the paper.
@@ II Related work
II-A Vascular ageing estimation methods
The works on vascular age estimation could be categorized into the following broad categories: statistical methods, machine learning methods, deep learning methods, and classical score based methods. In terms of hardware equipment, the non-invasive methods reported in the literature either utilize the 12-lead ECG or PPG sensors. For example, Ladejobi et. al. [5] provide a 12-lead ECG method for vascular age estimation. Similarly, Charlton et. al. [9] provide an in-depth and comprehensive review of the works that utilize PPG sensors to compute various parameters for the sake of vascular age estimation. Next, we consider selected works, group together the similar works under an umbrella term and discuss them below, one by one.
@@ II-A Vascular ageing estimation methods
Statistical methods: Ball et. al. [10] develop a Bayesian statistical model that predicts an individual’s heart age based on his/her resting 5 min 12-lead ECG. They evaluated their model on 776 healthy individuals, 221 with cardiovascular risk factors, 441 with cardiac disease, and a small group of highly endurance-trained athletes. Their model predicts a higher heart age for the subjects with risk factors, with heart diseases, and athletes. Motivated by the situations where cfPWV cannot be measured, Heffernan et. al. [11] compute the estimated pulse wave velocity (ePWV) from the age and brachial blood pressure. They measure cfPWV in two-hundred and fifty-two adults, and compare it with each participant’s ePWV, in order to correlate the ePWV with cfPWV (which is an established measure of vascular aging).
@@
Machine learning based methods: Recently, machine learning methods for vascular ageing prediction have gained widespread attention (see the survey article [7] and references therein for a quick review of the development in the field). Due to space constraints, we discuss selected works below. Kakadiaris et. al. [12] utilizes a 13-years MESA dataset of 6459 atherosclerotic CVD-free participants and demonstrates that their proposed support vector machine (SVM)-based CVD risk calculator outperforms the classical PCE-based CVD risk calculator. Dall Olio et. al. [13] utilizes a public dataset of video PPG signals acquired from smartphones, does the feature extraction, and trains a number of machine and deep learning methods in order to predict healthy vascular ageing. Suri et. al. [14] take a critical approach and study the bias problem in ML-based vascular age prediction methods. They argue that overfitting and lack of large datasets (with diverse ethnic populations) might be the reasons behind the fact that ML methods yield inaccurate results during the clinical trials. Similarly, Lindow et. al. [15] argues that it is possible to accurately predict patients’ Bayesian 5-min ECG Heart ages from their standard, resting 10-s 12-lead ECGs using standard regression models only, without relying upon the AI techniques. They utilize a public dataset with a total of 2,771 subjects (1682 healthy volunteers, 305 with cardiovascular risk factors, 784 with cardiovascular disease).
@@
Deep learning based methods: In [16], Lima et. al. train a residual neural network on the 12-lead ECG data from the CODE study cohort (1,558,415 patients), and report that the patients with a heart age gap of more than 8 years have a higher mortality rate. Hirota et. al. [17] train a convolutional neural network (CNN) on 12-lead ECG data of 17,042 subjects, and reports that the age predicted by their model matches well with the chronological age of young population, but not for the elderly (with age $\geq 60$ years). Chang et. al. [18] train the ECG12Net model (basically, a CNN with 82 convolutional layers) on 71,741 subjects (20 to 80 years old), and reports that the high ECG-age is not only correlated with mortality but with several CVD-related outcomes as well. In [19], Park et. al. collect PPG data from 757 participants, utilize a Gaussian mixture model to decompose the PPG signal into incident and reflected waves for feature extraction, and eventually train a shallow artificial neural network for vascular ageing prediction. Libiseller-Egger et. al. [20] train a CNN on 10-second 12-lead ECG data and CVD-related metadata from 499,727 subjects, with the aim to learn the genetic basis of cardiovascular ageing, in order to support future personalised/precision medicine (when genomic information is readily available). In [21], Attia et. al. train their CNNs using 10-second samples of 12-lead ECG signals from 499,727 patients in order to predict their biological age and gender. Benavente et. al. [22] train a CNN on over 700,000 ECGs from the Mayo Clinic, US, in order to later predict the ECG-age of 4,542 participants in two cities in Russia. Authors report that $\delta$-age (i.e., ECG-derived artificial intelligence-estimated age minus chronological age) biomarker is strongly correlated with the CVD risk factors, e.g., blood pressure, BMI, total cholesterol and smoking. In [23], Toya et. al. train a CNN on the 12-lead ECG data of 774,783 unique subjects, and conclude that the patients with peripheral microvascular endothelial dysfunction (an index of vascular aging) and accelerated physiologic aging are at a greater risk for CVDs.
@@
Low-cost single-lead ECG module design: Since this work utilizes the AD8232 chip (a low-cost single-lead ECG sensor with a price tag of 5 USD), it is imperative to summarize the works that design low-cost single-lead ECG module with the unanimous aim to provide affordable sensing solutions for remote health monitoring. There are works that consider a AD8232 chip-based low-cost single-lead ECG module for atrial fibrillation detection (a common cardiac arrhythmia associated with stroke risk) [27],[28]. Gromer et. al. [29] present a cost-effective method for detecting driver drowsiness, utilizing heart rate variability data from the ECG sensor. Further, works do exist that focus on the design of an internet-of-things (IoT)-based low-cost ECG system which reports various biomarkers (e.g., pulse rate, SpO2, body temperature, etc) of remote Covid-19 patients to the doctor, and could detect various cardiac abnormalities such as premature ventricular contraction, premature atrial complex, etc [30, 31, 32]. Porumb et. al. [33] aim to estimate blood glucose level in a subject using a non-invasive wearable ECG device, using features such as T-wave, ST-segment, etc.
Last but not the least, there are works that aim to estimate the biological age of a person using ECG or PPG data. The interested reader is referred to [34] and the references therein.
II-B Effect of smoking on ECG
According to epidemiological studies, Tobacco cigarette smoking is one of the (preventable) major risk factor for cardiovascular diseases, e.g., coronary artery disease, arrhythmia, atherosclerosis, ischemic heart diseases, and sudden cardiac death. This is because Nicotine, an ingredient of cigarette, alters the physiology of the heart by inducing conduction block, and by producing adverse effects on ventricular repolarization. The Nicotine also acts as a potassium channel blocker, and leads to abnormal ventricular repolarization, atrial depolarization, and atrioventricular depolarization, which in turn implies cardiac arrhythmia and more. Below, we describe selected works that have studied the impact of smoking on cardiac health through ECG waveform.
In [35], Yadav et. al. collect ECG data from male smoker and non-smoker subjects in the age range of 18-70 years, and observe a decrease in PR interval, decrease in QRS complex duration, change in R-wave amplitude and S-wave amplitude, change in ST segment duration, decrease in TP interval, and change in T-wave duration, for heavy smokers. Devi. et. al. [36] collect ECG data from 88 healthy male subjects (44 smokers, according to the ICD-10 criterion) in the age group of 18-30 years, in a hospital setting. They observe a decrease in QTc interval, increase in the QRS complex duration, increase in the heart rate, and decrease in the RR interval, the QT interval and the ST segment, for the smokers. Kumar. et. al. [37] collect 12-lead ECG data from 148 male patients, and observe an increase in mean systolic and diastolic blood pressure, prolonged QTc interval, increase in pulse, QTi, and QTc, and low mean respiratory rate interval, among smokers. In [38], Irfan et. al. analyze ECG data from 5,633 patients and observe a decrease in PR interval, PR segment, and QRS duration, for smokers. In [39], Ammar et. al., obtain 12-lead ECG from 105 subjects, with 35 being chronic e-cigarette users, and another 35 being conventional cigarette smokers. They observe an increase in mean heart rate, decrease in QRS complex duration, prolonged QT and QTc intervals, prolonged ventricular repolarization indices (T-wave-peak to T-wave-end (Tp-e) interval, Tp-e/QT ratio, and Tp-e/QTc ratio, among chronic e-cigarette users and conventional smokers. Gupta et. al. [40] performs 24-hours Holter monitoring on 60 male smokers and tobacco chewers (30 in each group) experiencing atypical chest pain, and observe an increase in the heart rate in both smokers and tobacco chewers, and an increase in supraventricular ectopics and a decrease in heart rate variability for smokers. Finally, [41] obtains ECG recordings from 37 chronic tobacco cigarette smokers and 43 chronic e-cigarette users, and compute Tp-e, Tp-e/QT ratio, and Tp-e/QTc ratio, in order to quantify the adverse effects of Nicotine on ventricular repolarization.
@@ III-A The Proposed Framework
Fig. 1 provides a detailed overview of our proposed method. As can be seen, we utilize our custom low-cost single-lead ECG sensor module in order to construct a labelled dataset. We then do pre-processing of the noisy ECG data, followed by segmentation, followed by feature extraction. Eventually, we pass a total of 21 features (13 ECG-based features and 8 other features) to a number of ML algorithms, a 1D CNN model, and the ResNet18 model which do vascular ageing prediction. We also do transfer learning whereby we pre-train our models on a public PPG dataset, and later, fine-tune and evaluate them on our unsegmented ECG dataset. In addition, we utilize the extracted features to identify smoking-induced changes in ECG.
Figure 1: The proposed method for vascular ageing prediction and for studying smoking-induced changes in single-lead ECG.
@@ III-B The Low-cost Single-Lead ECG Module
We utilize the AD8232 chip which is a low-cost single-lead ECG sensor (with a price tag of 5 USD). This ECG sensor requires a DC voltage in the range of 2-3.5 volts, and comes with three electrodes which enable it to capture the tiny voltages generated by the human heart, at a rate of 100 Hz. For systematic data acquisition, the AD8232 ECG sensor is interfaced with the ESP32-WROOM micro-controller board (developer version). Further, a 3.7 volts lithium-ion battery is used to power the ESP32 board, the ECG sensor, and TP4056 module (basically, a battery management system). This way, we realize a very cost-effective single-lead ECG sensor module (with a price tag of 10 USD). We connect and solder all the aforementioned hardware components on a veroboard, and then place it in an acrylic body designed via the Solidworks tool. We program the ESP32-WROOM and AD8232 ECG sensor using the Arduino IDE with the objective to read and log the ECG waveform for the purpose of our custom dataset construction.
@@ IV-A Data Collection
We utilized our custom module to collect single-lead ECG data from 42 apparently healthy subjects (26 male, 16 female) of age 18 years to 30 years, on campus [^1] This work was approved by the ethical institutional review board (EIRB) of Information Technology University (ITU), Lahore, Pakistan.. Each subject contributed to 3 minutes worth of data, leading to a dataset having 126 minutes (7560 seconds) of single-lead ECG data. Among the male subjects, 16 were smokers, while 10 were non-smokers. Similarly, 4 female subjects were smokers, while 12 were non-smokers. All the smoker subjects in this study are light but habitual smokers, i.e., they smoke only 1-2 cigarettes per day, but have been doing it on a regular basis since last 2-3 years. For the ground truth for the vascular ageing prediction problem, we utilized chronological age as a proxy for the vascular age (which is a reasonable assumption keeping in mind that we collected data from apparently healthy young subjects only) [15],[13]. In addition to acquiring the ECG traces of the subjects, we also recorded the relevant metadata such as age, height, weight, heart disease history in the family, smoking habits, sleep duration, blood pressure (systolic and diastolic), body mass index (BMI), and the resting heart rate.